OpenStudy (anonymous):

Courier charges for packages to a certain destination are 65 cents for the first 250 grams and 10 cents for each additional 100 grams or part thereof. What could be the weight in grams of a package for which the charge is $1.55 ? A. 1155 B. 1145 C. 1040 D. 950 E. 259

OpenStudy (anonymous):

Are you sure you've got the right answer choices?

OpenStudy (anonymous):

I keep getting 1150

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

100 grams or part thereof??

OpenStudy (anonymous):

let X be the weight in grams of package so 65 + (X-250)/100 x10 = 155 ==> 90 = (X-250)/10 ==> X= 1150 gm so we will have to pay $1.55 for 1150gm though $1.55 is valid for 1050 < X < 1150 So B is the ans..